Σ(-1)-F2-1 1=1

Guess formula and prove by induction
Answer #1

Here we prove que66 formula by voing induction ! iT where ai = c-vji- faian (Bay) for iar , 9 = (-1)Faci -= 4°F9-1 = Fi ize4= n then an = (-1,9-Fanal if nio odd then an =(-1) 19 Fendi = (I) Fen-1 o is even then an = (-1)^-! Fizn-J = -fan-1 Note: iby words induction suppose for n=1 them a = 1 = FI 1:2 for n=2 , then E i = 1 +92 - Fi - FS n:9 for ne then WE aj - +9nt 98 =